OpenStudy (anonymous):

Check my work: Inflection points Note: I think I may have found the possible inflection points. How do I check to make sure they are the right points? I know that they are this time, but the points are right next to each other.

OpenStudy (anonymous):

Inflection : f''(x) = 0

OpenStudy (anonymous):

I took the second derivative, and set it to 0. I got -1,0,and 1.

OpenStudy (anonymous):

So my answer is 3.

OpenStudy (anonymous):

60 x^3 - 60x = 0 ryt ?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

x^2 = 1

OpenStudy (anonymous):

x = 1 or -1

OpenStudy (anonymous):

Couldn't 0 also be a root?

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

is 0^2 = 1 ???

OpenStudy (hba):

No lol.